OpenStudy (anonymous):

write an equation for this problem? a number is 8 more than 32. find thw number.let n=number.

OpenStudy (ivettef365):

you know that 8 more than 32 = 32 + 8 and the number = n so what will be the equation

OpenStudy (anonymous):

n=32 +8

OpenStudy (anonymous):

thats what i thought the answer was but it said it was wrong....but thank u guys for helping me..

OpenStudy (ivettef365):

correct